Solving a linear equation with several occurences of the veriable problem type 2
5x+7/3=1/3x-7/2. How can I solve this question

Answer by stanbon(75887)

5x+(7/3) = (1/3)x-(7/2)
Multiply thru by 6 to get:
---------------------------
30x + 14 = 2x - 21
28x = -35
x = -35/28
===============
Cheers,
Stan H.